OpenStudy (anonymous):

A guy wire runs from the ground to the top of a 25 foot telephone. The angle formed between the wire and the ground is a 52 degrees. How far from tbe base of the pole is the wire to the ground?

OpenStudy (anonymous):

Can you draw a picture of this (not on here but on your paper)?

OpenStudy (anonymous):

there no pictue on my textbook sorry.

OpenStudy (anonymous):

I know, I am just wondering if you understand the picture?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

I will try and draw this top of tp | | \ | \ | \ | \ | \ | \ | \ | \ | \ | 52 \ ------------ The left hand side is 25 ft tall You are trying to find the bottom width So you know an angle which is 52 degrees You know the opposite side which is 25 ft You are trying to find the bottom which is the adjacent side which is x Tangent has to do with opposite over adjacent So tan(50) = 25/x tan(50) = 1.19 so 1.19 = 25/x multiply both side by x 1.19x = 25 divide by 1.19 x = ~21 ft.

OpenStudy (anonymous):

sorry to interject, but you could possibly draw the picture on paper, take a picture of it, and use the "Attach File" button to submit it here

OpenStudy (anonymous):

Don't have a scanner at home.... Thanks for the info tho..

OpenStudy (anonymous):

why is 52 degrees in the right side instead of the left

OpenStudy (anonymous):

It is the angle that joins the wire and the ground... The wire is the diagonal line.. between it and the ground is the 52 degrees